OpenStudy (anonymous):

Kingston and Dover are 150 miles apart. One train leaves Kingston traveling towards Dover at an average speed of 60 mph. Another train leaves Dover at the same time traveling towards Kingston at an average speed of 65 mph. a. How long will it take them to meet? b.How far has each train traveled when they meet?

OpenStudy (anonymous):

@zepdrix

OpenStudy (mathmath333):

u need to find the relative speed of both trains \(\large\tt \begin{align} \color{black}{relative~~speed\\ =s_1+s_2\\ =60+65\\ =125\\ distance =\dfrac{speed }{time }\\ time =\dfrac{125 }{150 }\\ time =\dfrac{5 }{6 }~~hrs\\ time =\dfrac{5 }{6 }\times 60~~min\\ \huge time =50~~min\\ }\end{align}\) distance travelled by train from Kingston when they meet \(\large\tt \begin{align} \color{black}{(d_1)distance =\dfrac{60 }{\dfrac{5 }{6 } } =\dfrac{60\times 6 }{5}\\ \huge d_1=72~miles}\end{align}\) distance travelled by train from Dover when they meet \(\large\tt \begin{align} \color{black}{(d_2)distance =\dfrac{65 }{\dfrac{5 }{6 } } =\dfrac{65\times 6 }{5}\\ \huge d_2=78~miles}\end{align}\)